OpenStudy (anonymous):

Help Please Factor completely: 16x4 – 1 A. (2x + 1)^3(2x – 1) B. (4x^2 + 1)(2x + 1)(2x – 1) C. (4x + 1)(4x – 1) D. (4x^2 + 1)(4x^2 – 1)

OpenStudy (anonymous):

16x^4-1

OpenStudy (anonymous):

u need that factored also

OpenStudy (anonymous):

(2x+1)^3(2x-1) Polynomial not factorable

OpenStudy (anonymous):

That's the original one, I just corrected it

OpenStudy (anonymous):

B. (4x^2+1) (2x+1) (2x-1) Answer: Not factorable C. (4x+1) (4x-1) Answer: Not Factorable D. (4x^2+1) (4x^2-1) Answer: (4x^2+1)(2x+1)(2x-1)

OpenStudy (anonymous):

To get the answer you're supposed to factor 16x^4-1, not factor the answer choices.

OpenStudy (anonymous):

oh lol im very sorry

OpenStudy (anonymous):

let me do that

OpenStudy (anonymous):

anyways the answer is D.

OpenStudy (anonymous):

Thanks for the medal

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

you can write this equation as \[((4x^2)^2 - 1^2) = (4x^2-1)(4x^2+1)\]

OpenStudy (anonymous):

by the formula a^2-b^2 = (a+b)(a-b)

OpenStudy (anonymous):

and there is more

OpenStudy (anonymous):

you see \[4x^2-1 \] can be written as \[(2x)^2-1^2 = (2x+1)(2x-1)\]

OpenStudy (anonymous):

so the final answer you get is \[(4x^2+1) (2x+1)(2x-1)\]

OpenStudy (anonymous):

which is option B :)

OpenStudy (anonymous):

easy,right?

OpenStudy (anonymous):

ohhhh im i see what u did

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

welcome